Please help!
If log b < 0, what can you say about b?
Thanks so much!

Answer by Theo(2978)

Let x = a positive number.

Then -x = a negative number.

log(b) < 0 implies that log(b) = -x because -x < 0.

Now log(b) = -x if and only if 10^(-x) = b.

This is by basic definition of logarithms.

10^-x is the same as 1/10^x by definition.

Since x >= 0 by definition, then the smallest 10^x could be would be 1 because 10^0 = 1.

Any other value of x > 0 would result in 10^x being greater than 1.

Example:

10^0.1 = 1.2589.....
10^0.00001 = 1.0000023026

Bottom Line is the smallest 10^x can be is 1.

Now, if 1/10^x = b, this means that the largest b can be is 1 because 1/1 = 1.

so, to answer your question:

If log(b) < 0, this means that 0 < b < 1

Some examples:

log(2) = .3...
log(1) = 0
log(.9) = -.04...
log(.5) = -.301...
log(.1) = -1
log(0) = Error - can only take log of a number > 0

So that's the answer to your question.

log(b) < 0 if and only if b is greater than 0 and b is smaller than 1.